10. An proton-antiproton pair is produced by a 2.20 x 100 MeV photon. What is the kinetic energy of the antiproton if the kinhi please show your work clearly and explain, thank you!
Answer #1

Energy of photon = Energy converted to mass + kinetic energy of anti-proton + kinetic energy of proton

2.2*10^3 MeV = 2*m*c^2 + 161.90 MeV + KE_proton

2.2*10^3*10^6*1.6*10^-19 = 2*1.67*10^-27*(3*10^8)^2 + 161.90*10^6*1.6*10^-19 + KE_proton

3.52*10^-10 = 3.265*10^-10 + KE_proton

==> KE_proton = 3.52*10^-10 - 3.265*10^-10

= 2.55*10^-11 J

= 2.55*10^-11/(1.6*10^-19) eV

= 1.5938*10^8 eV

= 159.38 MeV <<<<<<<<<<------------Answer